# Specialized Language Models with Cheap Inference from Limited Domain Data

### v1

@@ Abstract
Large language models have emerged as a versatile tool but are challenging to apply to tasks lacking large inference budgets and large in-domain training sets. This work formalizes these constraints and distinguishes four important variables: the pretraining budget (for training before the target domain is known), the specialization budget (for training after the target domain is known), the inference budget, and the in-domain training set size. Across these settings, we compare different approaches from the machine learning literature. Limited by inference cost, we find better alternatives to the standard practice of training very large vanilla transformer models. In particular, we show that hyper-networks and mixture of experts have better perplexity for large pretraining budgets, while small models trained on importance sampled datasets are attractive for large specialization budgets.
@@ 1 Introduction
Training large language models enables versatile models, but their high inference cost limits them to high-value applications Brown et al. (2020); Bommasani et al. (2022). Despite progress in approximated inference Aminabadi et al. (2022); Sheng et al. (2023); Dettmers & Zettlemoyer (2023), large models remain costly, or even impractical for mobile hardware. Under tight inference constraints, one might consider a small model specialized to the domain at hand. This paper studies training small specialized models, even with limited domain data. To achieve low perplexity, we use three key elements: generic training corpora, importance sampling, and asymmetric models with fewer parameters at inference than during training, such as mixtures of experts or hyper-networks.
With inference cost and in-domain training data limits, we study alternative strategies with varying training cost. We also take into account how training cost can be shared across domains. For our study, we consider 4 important metrics:
Generic training cost: the cost of the training phase that can be performed before the specialization data are available, on a generic training set. This cost is shared across multiple specialized models and is often called pretraining. Although not mandatory, generic training data are essential when specialization data are limited.
Specialization training cost: the cost of the training performed once the specialization data are available. This cost is not shared across different specialized models.
Inference cost: the cost for running inference on a specialized model. As part of the inference cost, one might also be interested in a model which involves a small number of parameters per specialized task, e.g. considering memory and network constraints. Low inference cost allows wider model deployment.
Size of the specialization training set: varies across applications and influences pretraining and specialization choices.
We take the inference cost and the specialization data size as hard constraints and study the operating curves resulting from varying the generic and specialization training costs. We compare different training approaches and highlight at which operating point they are interesting.
{forest} baseline [Inference budget?,l sep=8mm [Large Model,edge label=node[midway,left,font=]Large    ,draw] [Specialization budget?,l sep=8mm,edge label=node[midway,right,font=]    Small [Importance Sampling,edge label=node[midway,left,font=]Large    , align=center, base=bottom,draw] [Hyper-networks & Mixtures,edge label=node[midway,right,font=]    Small, align=center, base=bottom,draw] ] ]
@@
Figure 1: Practical recommendations for training language models that fit a predefined computational budget.
@@ 2 Methods
We consider different architectures to satisfy our inference constraint while leveraging a large generic pretraining set. We assess the drop in perplexity from our inference constraint compared to a larger model trained on the same data. Our recommendations are summarized in Figure 1.
2.1 Large Model
Large Model (LLM) One trains a large language model (LLM) on the generic data and uses this model as-is at inference Brown et al. (2020); Kaplan et al. (2020); Hoffmann et al. (2022). This approach requires a high pretraining cost but does not require specialization data and has no specialization cost. Inference cost for this method is high. Having never seen specialized data, it might be inaccurate when the specialization distribution is far from the pretraining distribution.
After generic pretraining, fine-tuning over specialization data can adapt the model Howard & Ruder (2018); Aghajanyan et al. (2021). This step usually improves perplexity but adds a specialization cost. This cost is limited when the amount of specialization data is small: early stopping limits fine-tuning to a few updates to avoid overfitting. Fine-tuning leaves the inference cost unchanged.
Parameter Efficient TuningOne fine-tunes only a subset of the parameters once the specialization data are available Hu et al. (2021); Lester et al. (2021); Houlsby et al. (2019). This strategy is advantageous if the specialization data are scarce because it mitigates overfitting. However, tuning fewer parameters may require more fine-tuning steps and hence increase the specialization cost. This method is practical when one needs to communicate only small model deltas for each new specialization. See Section F in the Appendix.
2.2 Small Model
Small Model (SLM) One trains a single small language model (SLM) before the specialization data are available and uses this model as-is at inference. This method does not require specialization data or incur any specialization cost. The inference cost for this method is low and so is the pretraining cost. However, a small model cannot use a large amount of generic data as well as a large model resulting in worse downstream performance. Similarly to larger models, fine-tuning can improve performance with an additional specialization cost.
No Pretraining (SLM-nopt) This method only trains the model on the specialization data. This is advantageous when the specialization budget and amount of specialized data are large or when the generic training distribution is very far from the specialization domain.
Importance Sampling (SLM-is) This method does not pretrain the model before the specialization data are available. Once the specialization set is given, SLM-is samples a tailored training set from the generic pretraining data to match the specialization distribution Xie et al. (2023). This method is a case of data selection Moore & Lewis (2010); Grangier & Iter (2022) and is advantageous when the specialization data are scarce. This method incurs a high specialization cost since pretraining on a (possibly large) tailored training set is necessary for each specialization domain. After pretraining, the model can be further fine-tuned over the specialization data.
Distillation (SLM-d) This method Hinton et al. (2015); Hsieh et al. (2023b); Zhu et al. (2023) uses a fine-tuned large model as a teacher and distills it in a small student model. Compared to the large teacher model, inference cost is lower, but accuracy is also reduced. Compared to a small model without distillation, the accuracy can be better. The teacher model provides rich targets from a model with better generalization, often reducing overfitting to the specialization set. Compared to SLM fine-tuning, this method has a higher generic pretraining cost: it requires generic training of a large model and a small model. Its specialization cost is also greater since it requires tuning the teacher model and then collecting the teacher outputs during distillation.
2.3 Hard Mixture of Experts (SLM-mix)
This method Eigen et al. (2014); Gross et al. (2017) divides the large pretraining set in smaller subsets, e.g. via clustering, and pretrains a small model (expert) on each part. Its pretraining cost and its overall number of parameters are high as both scale linearly with the number of clusters. Inference with a hard mixture is typically performed by forwarding each example to the expert corresponding to the cluster the example belongs to.
Once the specialization data are available, we specialize the mixture by selecting a single expert for each specialization task. One option is to select the expert whose pretraining cluster is the most frequent cluster in the specialization data. Alternatively, if the specialization budget is sufficient, we can select the expert with the smallest loss on average on the specialization data. Selecting a single expert per domain is advantageous as it communicates and loads only a small part of the mixture weights for inference on the target domain, but it might be detrimental when the specialization data are spread across multiple clusters.
SLM-mix can be fine-tuned. With a large specialization budget, one can fine-tune each expert and select the best-performing one. With a smaller budget, one can instead fine-tune just the best-performing expert at pretraining or the expert whose pretraining cluster is the most frequent cluster in the specialization data. This second option makes the cost of specialization identical to SLM fine-tuning.
2.4 Hyper-Networks (SLM-hn)
Hyper-networks Ha et al. (2017) are neural networks that decompose into two parts: the hyper-sub-network and the instantiated sub-network. The hyper-sub-network creates weights for the instantiated sub-network. We rely on hyper-networks to create a mixture of experts: the hyper-sub-network takes the cluster membership of an input to produce the sub-network weights. These cluster-specific weights instantiate a small sub-network or expert. Compared to a hard mixture of experts, SLM-hn shares parameters across experts via the hyper-sub-network and provides a flexible way to independently select the capacity of the mixture and the number of clusters. An instantiated sub-network can be fine-tuned on specialization data.
@@ 3 Experimental Setup
We present the datasets for our experiments, the experimental setting for each method and the evaluation metrics.
@@ 3.1 Datasets
Our generic pretraining set is c4, a large filtered dataset of English text derived from commoncrawl Raffel et al. (2020). We tokenize the data with a sentence piece model trained on c4 with a vocabulary size of 32k. We consider specializing to nine diverse domains, extracted from the Pile Gao et al. (2021): arxiv (science articles), europarl (parliamentary proceedings), freelaw (legal text), gutenberg (old books pusblished before 1919), opensubtitles (theatrical subtitles), openwebtext2 (forum discussions), pubmed-abstracts (medical article abstracts), stackexchange (Q&A mostly about technical topics), wikipedia (encyclopedia articles). We vary the amount of specialization training data available and consider sets of size 1, 8 and 64 million tokens for each domain.
3.2 Clustering
Hard mixture-of-experts, hyper-networks and importance sampling rely on document clustering. We use sentence BERT Reimers & Gurevych (2019) to embed the c4 documents into 768-dimensional vectors and cluster them with the kmeans algorithm. We explore different numbers of clusters ranging for 4 to 1,024 clusters.
3.3 Language Models
We perform our experiments with transformer models Vaswani et al. (2017). We consider two model sizes, small and large. The small model has 126M parameters and consists of 7 layers with a dimension of 1,024 and a latent feed-forward dimension of 4,096. Our large model has 770M parameters with 7 layers with a dimension of 2,816 and a latent feed-forward dimension of 11,264. Models are trained and evaluated with a context of at most 1,024 tokens, splitting longer documents into non-overlapping windows.
3.4 Distillation
For distillation, we use a fine-tuned LLM as the teacher and an SLM pretrained on the generic set as the student. Distillation training operates on the specialization data and trains the student to minimize the KL divergence between its prediction and the teaching distribution, a mixture between the data distribution and the teacher model prediction Hinton et al. (2015). The teaching mixture weight is a hyperparameter (0.95 in our experiments). In this method, the generic training cost is dominated by the training of the teacher model while the specialization cost is also dominated by the cost of fine-tuning the teacher model. This method has an additional smaller cost of pretraining the SLM on the generic dataset and teaching the SLM on the specialization dataset.
3.5 Mixture of Experts
We train hard mixtures Gross et al. (2017) of transformers for pretraining. The pretraining set is divided into clusters and an independent SLM is trained on each cluster. For specialization, we consider a simple strategy: we cluster the specialization set with the pretraining centroids to determine the most frequent cluster in the specialization set. We fine-tune only the model pretrained on this cluster. Hard mixtures are interesting here since they allow training a model with a large total number of parameters while fine-tuning and running inference only with a small model.
If the pretraining budget is low, one can forgo pretraining models on all clusters and, instead, increase the specialization budget to train a model only on the cluster of the generic dataset corresponding to the most frequent cluster in the specialization set, once this set is available.
3.6 Hyper-Networks
Hyper-network Ha et al. (2017); Karimi Mahabadi et al. (2021) defines the general idea of a neural network whose weights are themselves generated from a secondary network, the hyper-network, based on a conditioning input variable. In our case we associate each example with its cluster membership variable, using the clustering mentioned in Section 3.2. This variable is the input of the hyper-network which produces the feed-forward (i.e. multi-layer perceptron, MLP) matrices of a transformer language model for all layers except the first two. The other parameters of the transformer do not depend on the cluster and are the same for all examples.
Our hyper-network instantiates two MLP matrices $W^{(1,l,i)},W^{(2,l,i)}$ for each layer $l$ and each cluster $i$. It relies on two hyper-parameters: the latent dimension $h$ and the number of experts $m$. Each cluster $i$ is associated with the h-dimensional embedding $c^{(i)}$. Each layer l is associated with the $h\times m$-matrix $M^{(l)}$. We compute the matrices
$W^{(1,l,i)}=c^{(i)}~{}M^{(l)}\cdot T^{(1,l)}\textrm{~{}and~{}}W^{(2,l,i)}=c^{(i)}~{}M^{(l)}\cdot T^{(2,l)}$
as the weighted sum between the vector $c^{(i)}~{}M^{(l)}$ and the three dimensional tensors $T^{(1,l)},T^{(2,l)}$ respectively of shape $m\times d_{\text{latent}}\times d_{\text{in}}$ and $m\times d_{\text{in}}\times d_{\text{latent}}$. These tensors hold most of the model parameters, i.e. $m$ times as many parameters as the corresponding MLP matrices. This strategy enables increasing $m$ to increase the overall model capacity while keeping the size of the model instantiated for each cluster constant. Of course, we illustrate one choice of hyper-network architecture, but many alternatives are possible Muqeeth et al. (2023); Abnar et al. (2023). Compared to hard mixtures of experts, the hyper-networks have stronger capacity limitations since the training problem cannot be split into independent, low-memory training tasks. On the other hand, the weights of each of the $m$ experts — both the attention parameters, which are the same for all experts, and the MLP tensors — are trained jointly, hence the hyper-network model can be more parameter efficient.
For specialization, we follow a strategy similar to the hard mixture case: we instantiate the model at the most frequent cluster on the specialization set and fine-tune it. Fine-tuning therefore does not operate on the large hyper-network but only on the small instantiated model.

### v2

# Need a Small Specialized Language Model? Plan Early!

@@ Abstract
Large language models are versatile tools but are not suitable for small inference budgets. Small models have more efficient inference, but their lower capacity means that their performance can be good only if one limits their scope to a specialized domain. This paper explores how to get good specialized small language models using a large, generic, pretraining set and a limited amount of specialized data. We consider two scenarios, depending on whether (i) one can afford pretraining a model for each specialization task, or (ii) one wants to cheaply adapt a single pretrained model for each task. In the first scenario, we propose an effective solution based on importance sampling: we resample the pretraining set to imitate the specialization data and train a small model on it. In the second scenario, we propose a novel architecture, projected networks (PN). PN is a large network whose parameters can be linearly projected into a small network for specialization. For both scenarios, we demonstrate the empirical effectiveness of our solutions across various domains, training set sizes, and training budgets.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a generic tool to address a wide range of language tasks (Brown et al., 2020; Bommasani et al., 2022). This generality requires a large, diverse, generic training set. This rich set ensures that the model fits many subdomains close to the tasks it will eventually address. Model generality is particularly impactful for tasks where the cost of collecting a representative training set cannot be justified. However, LLMs inference is costly because of their large number of parameters, required to fit a large training set. This cost restricts LLMs to high-value applications. Efficient inference is an active research area that follows multiple routes like model distillation (Hsieh et al., 2023), quantization (Dettmers & Zettlemoyer, 2023), pruning (Ma et al., 2023) or hardware optimization (Aminabadi et al., 2022). However, reducing model size is the most direct solution for applications under tight inference constraints, and it can be combined with all the aforementioned techniques to further reduce inference costs. With that inference constraint in mind, we focus on training Small Language Models (SLMs).
A SLM cannot fit a generic training set as well as an LLM (Vapnik, 1995; Bishop & Bishop, 2023). It is hence necessary to forgo the generality of the model to devote its limited capacity to a targeted specialization domain. While a large specialized training set for the application at hand would be ideal for training, such a set is costly and usually justified only for high-value applications. Many applications, therefore, have to face both a limited inference budget and a limited in-domain training set size. For instance, some applications cannot afford to collect more than a few million tokens worth of training data (1m tokens $\simeq$ 10 books). Such applications, with low inference and low data collection budgets, are facing a challenging problem.
This work tackles precisely this problem. We aim to answer the following question: How can we get a specialized small language model when little specialization data is available? To answer this question, we distinguish two scenarios of practical importance, each leading to different recommendations: (i) when one needs a single small model for one specialized domain, a large training budget can be considered for that domain and one can pretrain a model specifically targeted to this domain; on the other hand, (ii) when one needs one model per domain for many specialized domains, pretraining one model per domain quickly becomes too expensive, and it is interesting to share the training cost across domains.
For both scenarios, there are natural baselines that come to mind. The first baseline consists of pretraining a small model using the large generic pretraining set and then fine-tuning it with the little available specialization data. Another widely used method is distillation (Hinton et al., 2015): one pretrains a large teacher and a small student model on the pretraining set, fine-tunes the large model on the specialization set, and then fine-tunes the small model with the distillation loss from the fine-tuned large model.
While these two baselines are widely used in practice, for both scenarios, we propose improved pretraining strategies and demonstrate their superiority. The cornerstone of our improved strategies is a clustering of the pretraining set, which allows us to sample data points from each cluster of the pretraining set and pretrain on a mixture of these clusters with the flexibility to choose the weight of each cluster. When one targets a single domain (i), we find that an effective method is pretraining a small model over generic data with importance sampling (Owen, 2013). We pretrain on a mixture of the clusters where the weights are chosen to mimic the targeted domain. We show that, even though this emphasis relies on information from the specialization data, such a pretrained model still benefits from fine-tuning over the specialization data and outperforms the previous baselines. When one targets multiple domains (ii), we propose a novel architecture for pretraining, the projected network (PN). This model is a large model trained on generic data in which some parameters are tied to a cluster. Crucially, its parameters can be linearly projected into different small models — one per cluster — prior to fine-tuning. When one is given a new specialization domain, we show that it is effective to select a single projection and fine-tune the corresponding small model. In this paradigm, training the generic high-capacity model is costly, but its specialization phase (projection onto an SLM and SLM fine tuning) is not. This is ideal when many specialized models are needed.
We evaluate each method on several domains, pretraining and specialization budgets as well as specialized dataset sizes. Compared to the baselines, both importance sampling and projected networks bring a strong improvement, on all considered domains, specialized dataset sizes and training budgets. We show that importance sampling results in better specialized-perplexity than projected networks. However, importance sampling incurs a high pretraining cost when targeting multiple domains since pretraining is not shared across domains. As a result, projected networks are beneficial in the second scenario, for applications requiring many specialized models.
Contributions In Section 2, we present classical pretraining strategies as well as our novel methods based on a clustering of the pretraining dataset. We present our cluster-based importance sampling strategy. We introduce Projected Networks, a novel application for hard mixture models and hyper-networks, and explain how this provides multiple models that can be quickly instantiated during the specialization phase. In Section 3, we propose a methodology to evaluate different strategies to train a specialized SLM and identifies four important variables: the generic training budget (for training before the target domain is known), the specialization budget (for training after the target domain is known), the inference budget, and the in-domain training set size. In Section 4, we provide a comprehensive empirical study by reporting experiments on 9 domains, 3 specialized training set sizes with various training budgets. We show that, as expected, fine-tuning is essential. Surprisingly, we highlight that distillation from a large model, albeit popular, brings little improvement when accounting for the overall pretraining budget. We show that our cluster-based importance sampling method is very effective for training specialized models in the large specialization budget case. Finally, we find that Projected Networks lead to the best models in the small specialization budget case. Figure 1 summarizes our practical recommendations.
@@
Figure 1: Practical recommendations for training LMs that fit a predefined computational budget.
@@ 2 Methods
We consider different strategies to leverage a large, generic pretraining set under inference constraints.
2.1 Generic and Specialization Datasets
We have access to a large, generic pretraining set, which is typically obtained as the output of a web crawl. This set is large: it has enough samples to train a model on it or on a subset of it without risks of overfitting. This set is generic: it covers a wide variety of topics and concepts. We train models on this set using the next-token prediction loss.
We also have access to a small specialization set, which can be, for instance, the internal documentation of a company, some emails, or a set of scientific articles. This set is small: one cannot train a good model on it without overfitting. This set is specialized: it only covers specific topics. Once again, we measure the quality of models on this set with the next-token prediction loss and perplexity.
2.2 Baselines: Small Models, Fine-Tuning & Distillation
We target a model with a good specialized perplexity under inference constraints. We abstract the inference constraints as a limit on the model size. We call Small Language Model (SLM) a transformer of the maximal possible size allowed by the inference constraint.
We consider 3 variants of SLMs. We first consider an SLM model trained only on the generic pretraining data. This model can be effective if the generic and specialization data are close. We then consider an SLM trained only on the specialization data . This model can be effective if the specialization training set is large enough but will quickly overfit when the specialization set is small. Finally, we consider a model pretrained on the generic data and fine-tuned on the specialization data. The amount of fine-tuning can be adjusted via early stopping. This adjusts a trade-off between the proximity to the generic distribution and overfitting to the specialization set. Training and fine-tuning are done by minimizing the next-token prediction loss on the corresponding datasets.
Since our capacity constraint is motivated by inference efficiency, we can consider larger models at training time and rely on distillation (Hinton et al., 2015) to satisfy the inference requirements. Specifically, we consider a large language model (LLM) that has a bigger size than the inference constraint allows. We pretrain it on the generic pretraining set, and then fine-tune it on the specialization set, achieving a lower perplexity on the specialization set than the pretrained-then-fine-tuned SLM thanks to its larger size. A pretrained SLM student is then fine-tuned to a mixture of the specialization data and next-word distributions from the teacher. We call this model SLM-d.
This type of distillation is often called Response-based Knowledge distillation in the literature (Gou et al., 2021, Sec 2.1) and is arguably the most widely used type of distillation. We do not consider using the LLM as a data augmentation tool (Feng et al., 2021), to produce synthetic data (Tang et al., 2019; West et al., 2021) resembling the specialization data, and then use them to train or fine-tune the small model. Indeed, this requires training an LLM, which is time-consuming, and in our setup, this leads to more variance than response-based knowledge distillation and overall poorer performances (Menon et al., 2021).
2.3 Clustering of the Pretraining Data
The subsequent strategies rely on a clustering of the generic pretraining set. Clustering associates each training example with a discrete latent variable that we use to efficiently estimate importance sampling weights with a conditional independence assumption (Section 2.4). It also allows us to condition the projection in projected networks (Section 2.5).
To cluster the data, we embed each document in the generic set as a vector using Sentence BERT (Reimers & Gurevych, 2019). The documents longer than the maximum context we consider (1,024) are broken into nonoverlapping windows. Then, we use the k-means algorithm to cluster the generic set into $k$ clusters. We can then query samples from each cluster.
2.4 Cluster-Based Importance Sampling
Importance Sampling (IS) is a well established method (Owen, 2013) that enables estimating the expectation of a variable (the next-token prediction loss in our case) on a distribution (the specialization distribution), while sampling from a different distribution (the generic distribution). In practice, IS needs to estimate importance weights from data. We leverage the pre-training set clustering proposed above to make the importance weight estimation easy, by assuming our data are sampled from a mixture over clusters. The specialization loss is
${\cal L}(D_{\text{spec}};\theta)=\mathbb{E}_{x\sim D_{\text{spec}}}[\ell(x;\theta)]=\sum_{x}\ell(x;\theta)P(x|D_{\text{spec}}).$
We introduce a latent cluster membership variable $c$, and make an independence assumption $P(x|c,D_{\text{spec}})=P(x|c)$, which gives
${\cal L}(D_{\text{spec}};\theta)=\sum_{x}\sum_{c}\ell(x;\theta)P(x|c,D_{\text{spec}})P(c|D_{\text{spec}})=\sum_{x}\sum_{c}\ell(x;\theta)P(x|c)P(c|D_{\text{spec}}).$
We then apply importance sampling,
${\cal L}(D_{\text{spec}};\theta)=\sum_{x}\sum_{c}\ell(x;\theta)P(x|c)\frac{P(c|D_{\text{spec}})}{P(c|D_{\text{generic}})}P(c|D_{\text{generic}})=\mathbb{E}_{x\sim D_{\text{generic}}}[w(x)\ell(x;\theta)]$
with the importance weight $w(x)=\frac{P(c(x)|D_{\text{spec}})}{P(c(x)|D_{\text{generic}})}$ and $c(x)$ denotes the single cluster $c$ such that $P(x|c)>0$. The importance weights can, therefore, be estimated as the ratio between the cluster frequencies in the generic and specialization training set. The number of cluster $k$ is a trade-off between large $k$ (unreliable cluster frequency estimates for the specialization set, risk of overfitting to that set) and small $k$ (stronger independence assumption when the clusters are large). The small models trained with importance sampling are called SLM-is.
2.5 Asymmetric Models: Projected Networks and Hard Mixtures
Our inference constraints require that the final specialized model is a low-capacity SLM. Prior to fine-tuning on the specialization data, the capacity limit does not apply to generic pretraining. We devise a pretraining strategy to take advantage of this asymmetry. At pretraining time, we train a network with many parameters, but each example only interacts with a projection of the parameters onto an SLM. Like distillation, this strategy trains a model with many parameters, but unlike distillation, all model evaluations during training are already constrained to operate within the size limits.
[…]
$W^{(l,i)}_{a,b}=\sum_{q=1}^{m}E_{i,q}\sum_{r=1}^{h}M^{(l)}_{q,r}~{}T^{(1,l)}_{a,b,r}\text{ for }a=1\dots d\text{ and }b=1\dots d^{\prime}$
where $E_{i}\in^{m}$ is an expert-specific vector, $M^{(l)}\in^{m\times h}$ is a layer-specific matrix and $T^{(1,l)}\in^{d\times d^{\prime}\times h}$ is a tensor that stores most of the parameters. In our experiments, our SLM-pn experts are transformers and we only apply the PN decomposition to the feed-forward layers (i.e. multi-layer perceptron, MLP) which hold most of the model parameters. The other parameters are shared.
[…]
Figure 2: Projected networks (right) unlike distillation (left) instantiate small models in closed-form.
[…]
@@ 3 Experimental Setup
3.1 Methodology
With inference cost and specialization training data constraints, we study the alternative training methods at various training costs. We report training costs and which part of the cost can be shared across multiple domains. We consider 4 important metrics:
Generic training cost: the cost of the training phase that can be performed before the specialization data are available, on a generic training set. This cost is often called pretraining. It is domain-independent and can be shared across multiple specializations, e.g., via later fine-tuning. Although not mandatory, the generic training data are essential when specialization data are limited.
Specialization training cost: the cost of the training performed once the specialization data are available. This cost is not shared across different specializations.
Inference cost: the cost of running inference on a specialized model. Low inference cost allows wider, cheaper model deployment.
Size of the specialization training set: it varies across applications and influences pretraining and specialization choices.
Taking the inference cost and the specialization data size as hard constraints, we study the operating curves resulting from varying the generic and specialization training costs. We measure training cost (pretraining and specialization) in hours of graphic processor compute time (GPUh) on the same hardware (Nvidia-A100). We consider pretraining costs ranging from 10 to 650 GPUh and specialization costs ranging from 0.3 to 120 GPUh.
We evaluate language modeling with perplexity, using 20k held-out documents per dataset. We focus solely on language modeling; evaluating the models on downstream tasks (e.g. question answering, sentiment analysis, translation, etc) is beyond the scope of the paper. However, our conclusions could extend to downstream tasks as perplexity and downstream performance are often correlated Gonen et al. (2022); Du et al. (2024); Gadre et al. (2024). We report perplexity on generic and specialization data. For the different specialization domains. We report perplexity per domain in Appendix E and present macro-averaged results in Section 4. For macro-averaged perplexity, we compute the mean negative log-likelihood per token for each domain, average these results, and compute the exponential. All domains therefore get the same weight, regardless of the number of tokens per held-out set.
@@ 3.2 Datasets
Our generic pretraining set is c4, a large filtered dataset of English text derived from commoncrawl (Raffel et al., 2020). We tokenize the data with a sentence piece model trained on c4 with a vocabulary size of 32k. We use the clustering method described in Section 2.3 to split this dataset into $k$ clusters. We use $k=1024$ for SLM-is, and $k=32$ for SLM-pn and SLM-mix. We investigate the impact of the number of clusters in Appendix D.
@@
We consider specializing to nine diverse domains, extracted from the Pile (Gao et al., 2021): arxiv (science articles), europarl (parliamentary proceedings), freelaw (legal text), gutenberg (old books pusblished before 1919), opensubtitles (theatrical subtitles), openwebtext2 (forum discussions), pubmed-abstracts (medical article abstracts), stackexchange (Q&A mostly about technical topics), wikipedia (encyclopedia articles). We vary the amount of specialization training data available and consider sets of size 1, 8 and 64 million tokens for each domain.
3.3 Models Hyper-parameters
Table 2 reports the number of parameters for the pretrained and specialized models. Table 2 illustrates that SLM-pn and SLM-mix (Section 2.5) are as small as SLM for inference after specialization while their overall number of pretrained parameters is larger than LLM. Table 2 reports the throughput of the models. All SLM models have the same specialization throughput while SLM-pn has a lower throughput than SLM, SLM-mix for pretraining. LLM is more expensive in all cases. Table 3 presents the upper limit in training budgets for pretraining and specialization over all settings. Appendix A reports the training hyperparameters.
Table 1: Number of parameters (millions) for generic pretraining and inference. SLM-mix and SLM-pn are large models during pretraining but small at inference.